OpenStudy (anonymous):

can someone show me the first step in how to solve the trig equation: 2cos^2x-cosx-1=0. the restrictions are 0

OpenStudy (anonymous):

im not sure what identity to use here

OpenStudy (jhannybean):

is it \[\cos^2 x \] or \[\cos ^{2x}\]?

OpenStudy (anonymous):

cos\[\cos ^{2} x\]

OpenStudy (jhannybean):

\[2\cos ^{2}x-cosx-1\] okay. well do you know the identity for\[\cos ^{2}x\]?

OpenStudy (anonymous):

1+cos(2x)/2 ??

OpenStudy (jhannybean):

Yes :D so now you can replace it ! \[2(\frac{ 1+\cos(2x) }{ 2 })-cosx-1=0\]

OpenStudy (jhannybean):

Oh wait, i'm sorry,you don't even have to do this. this is much like the previous question you asked. Sorry sorry.

OpenStudy (anonymous):

haha im so confused on how to begin this because I have the same equation written down but I get totally stuck after that.

OpenStudy (jhannybean):

So lets start over. D: I jumped ahead of myself. You have \[2\cos ^{2}x-cosx-1=0\] you can write that in terms of x, right? \[2x^2-x-1=0\]

OpenStudy (anonymous):

yeah...

OpenStudy (jhannybean):

So factoring that our you have \[2x^2-x-1=0\] and this is a quadratic equatino solving for x! :) where you'll have 2 values for x. \[x^2-x-2=(x-\frac{ 2 }{ 2 })(x+\frac{ 1 }{ 2 })\]\[(x-1)(2x+1)\]

OpenStudy (jhannybean):

substitute your \[x=\cos x\] BACK into the equation and you'll get \[(\cos x-1)(2\cos x+1)\] and now we have to solve each equation separately. \[\cos x=1\] and \[2\cos x=-1 \rightarrow \cos x=-\frac{ 1 }{ 2 }\]

OpenStudy (jhannybean):

Do you see where i'm going with this?

OpenStudy (anonymous):

no not really I have been using a different method to solve them? haha

OpenStudy (jhannybean):

Oh.. which method have you been using to solve them?

OpenStudy (jhannybean):

Basically you're evaluating the equation to find x-values that correspond with the angle measurement. so if cos (x)=1, you're trying to find all the radian measurements where cos (x) will equal 1. and the same thing with cos (x) = -1/2

OpenStudy (anonymous):

okay I have been doing my whole workseet wrong because I didn't understand what I was trying to do. but that makes sence but then again I might not be doing it wrong. for example, one that I thought I got the answer to is \[\cos ^{2} x-3\sin x=3\] I worked it out like this : \[1-\sin ^{2} x-3sinx=3\] \[\sin ^{2} x+3 sinx+2\] \[sinx+1(sinx+2)\] =sin(x)=-1 -1=90degrees so x=90degrees

OpenStudy (jhannybean):

There are many ways to work the problem, but you are working it fine :) As long as you are finding the radian measurement of the angle in question.You got it. But did you equal both of those to zero? \[sin(x) +1 =0\] and \[\sin(x)+2 =0\]?

OpenStudy (anonymous):

whoops!(: I was thinking to far in advanced that's a mistake!

OpenStudy (jhannybean):

Mmhmm

OpenStudy (jhannybean):

Understand how to solve this problem now? :P

OpenStudy (anonymous):

yesss haha thank you!

OpenStudy (jhannybean):

No problem! ^_^